OpenStudy (anonymous):

Simplify 5 log2 k – 8 log2 m + 10 log2 n

OpenStudy (anonymous):

First of all use the property : \[\huge \color{red}{2Logx=Logx^2}\]

OpenStudy (anonymous):

Make sense ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Get the number to the log as a power As an example : \[\Large \color{blue}{5Log_{2}k=Log_{2}k^5}\] Can u do that to the rest ?

OpenStudy (anonymous):

ok so its log2k^5-log2m^8+log2n^10 but how do i add and subtract this.

OpenStudy (anonymous):

Great,So here is another property : \[\Large ==> \color{orange}{LogX+LogY=Log(X*Y)}\] \[\Large ==> \color{orange}{LogX-LogY=Log( \frac{x}{y})}\] If we collect both of them in one question As: \[\LARGE ==> \color{red}{LogX+LogY-LogZ-LogK= \frac{LogX*LogY}{LogZ*LogK}}\]

OpenStudy (anonymous):

ooooh i got it thanks

OpenStudy (anonymous):

Yw,Inform me when you need any help :).

OpenStudy (anonymous):

ok thanks again